OpenStudy (anonymous):

Use the definitions of the trigonometric ratios for a right triangle to derive a cofunction identity for tan(90° − A). Show your steps.

OpenStudy (anonymous):

is my answer correct tan(90° − A) = tan B .= b/a = cot A

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Since triangle ABC is a right triangle with C = 90, then A + B = 90, and B = 90 - A

OpenStudy (anonymous):

so did i answer correctly

OpenStudy (mathstudent55):

tan B = opp/adj = AC/BC but B = 90 - A, so tan (90 - A) = tan B cot A = adj/opp = AC/BC = tan B tan(90 - A) = cot A

OpenStudy (mathstudent55):

you got it right

OpenStudy (anonymous):

o thanks so much